1. If a gas under 5 atmospheres of presssure has a volume of 2 liters and a temperature of 10 degree celsius, What will be the new volume if the pressure increases to 7 atmospheres and the temperature rises to 50 degrees celsius?
A. 1.6 L b. 3.7 L C neither A nor B

2. 300 g of ice at zero degree centigrade is mixed with 500 g of water at 50 degrees centigrade. What will be the final temperature of the mixture? A. 1.25 celsius B. 2.0 Celsius C. neither A nor B

3. What is the change in lenght of a 5 m long metal rod if the metal has a coefficient of expansion of 4 X 10^-4/degree celsius and the temperature rises from 20 degrees centigrade to 120 degree centigrade? A. 0.2 m B. 0.008 m C. 0.53 cm
4. A rod having a crosssectional area of 0.3 cm^2 is hung vertically with a 1000 N weight hangiig on one end. If Young's Modulus is 4 X 10^8 N/m^2 and the original length is 4 meters,
What is the change in length when the weight is hanging on one end.
A. 3 cm B. 6.29 cm C. 0.3 m

5. A block is 5 cm X 7 cm X 2cm and has a mass of 128 g. If the density of water is 1 g/cc what is the apparant mass for the block once its submerged in Water? A. 58 g B. 70 g C. 140 g

6. Water is in a container that is open at the top and has hole at the bottom that water can escape. If the container is 2 meters deep and the container has a pressure of 50 Pa applied to the top surface, what is the flow rate from the opening at the bottom? A. 6.3 m/s B. 11.8 m/s C. 15 m/s

7.The first law of thermodynamics states heat added to a system results in work done and/or change in interanal energy of the system. A. True B. False

8. A metal rod has a length of 5 m and undergoes a temperature change of 50 degree celsius. If the coefficient of expansion is 20 X 10^-6/ degree celsius, what is the change in length of the rod? A. 500 mm B. 5 mm C. 6.3 ft

9. How much heat in calories is needed to change 30 g of ice at -7 degrees celsius to water at 40 degrees celsius?
A. 3000 cal B. 1214 cal. C. 3705 cal

10. What is the efficiency of a heat engine that has an input tmperature of 500 degrees celsius and an exhaust of 250 degrees celsius? A. 50% B. 32 % C. 22%.
Answer #1

2 = P = Satom V=22 Ti=10°c -> Ti = 10 +273 = 283.K P2 Tatm T2=500 V? T2 = 50+273 = 323 K we know 9. V = nRT mo. of moles R= G